An eight-point star-shaped polygon has a point angle of 36 degrees. What is the measure of its dent angle?. Also can you please show the work on how you came up with the solution?
Answer by Alan3354(69443)

An eight-point star-shaped polygon has a point angle of 36 degrees. What is the measure of its dent angle?. Also can you please show the work on how you came up with the solution?
-------------
8 points --> 16 sides
The total interior angles of a 16-gon = 180*(14) = 2520 degs
8 points * 36 = 288 degs
2520-288 = 2232 degs for the other 8 angles
2232/8 = 279 degs each